OpenStudy (anonymous):

Jack tries to define a function g: ℚ → ℚ by the rule g(m/n) = m - n for all integers m and n with n ≠ 0. Susan claims g is not well defined. Who is right? Justify your answer

OpenStudy (kinggeorge):

To not be well defined, you need to show that there is some (m/n) that gets mapped to two different values.

OpenStudy (kinggeorge):

At first glance this may seem well defined, but a number in \(\mathbb{Q}\) can be represented in several ways. For example, \[\frac{1}{2}=\frac{2}{4}=\frac{3}{6}\]If it were well defined, each of these values should have the same output.

OpenStudy (anonymous):

ohhhh i get it. sorry these r just discussion questions that i have never even briefed the topic so like there is no way that i can even give my input

OpenStudy (kinggeorge):

Don't worry about it :)

OpenStudy (anonymous):

Thanks gonna head to bed. I have had a rough day and my brain is closing shop. Thanks KingG U r one devoted fella that always comes to help me :)

OpenStudy (kinggeorge):

You're welcome.